OpenStudy (anonymous):

Trinh is catering a wedding that 112 people will attend. She is making two cakes, a chocolate cake and a carrot cake. of the guests will be served chocolate cake and the rest will receive carrot cake. How many guests will get carrot cake?

OpenStudy (anonymous):

You're missing a number in there. How many guests got chocolate?

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

alright, so 5/8*112 = nb of people getting chocolate = 70 people the number of people getting carrot : 112-70 = 42 people

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Eddie is fishing. He has 1 1/4 ounces of weight on his line, but his bait isn’t getting to the bottom of the lake. Eddie adds another 1/2 ounce weight to the line, but it's still not enough. Finally, Eddie adds a 3/4 ounce weight and the bait sinks. How much weight does Eddie now have on his line?

OpenStudy (anonymous):

You can try seeing it this way : in every group of 8 people, 5 will have gotten chocolate. First thing you do is find how many groups of 8 people you have (14 in this case). Then, in those 14 groups, you have 5 people getting chocolate, so 14 groups * 5people per group, that gives you 70

OpenStudy (anonymous):

ok, so you have 1 + 1/4 as initial weight 1 + 1/4 + 1/2 + 3/4 is the total weight at the end. you have to put everything on the same denominator to add these terms : 1 = 4/4 1/2 = 2/4 so your equation is now 4/4+1/4+2/4+3/4 = 10/4 10/4 is also 2 and a half ounces (depending on how your teacher wants the answer)

OpenStudy (anonymous):

Sam bought 2 1/4 pounds of Swiss cheese and 1 1/2 pounds of cheddar at the deli. How many pounds of cheese did Sam buy altogether? so do i just add them up?

OpenStudy (anonymous):

exactly, and how will you proceed with doing that?

OpenStudy (anonymous):

2 1/4 = 9/4? 1 1/2= 3/2 add and get 12/6

OpenStudy (anonymous):

or 12/4

OpenStudy (anonymous):

not exactly. I see what you did, and the only problem is you don't have the same denominator. Let's use coins instead, that may make it simpler to explain. So you have a quarter and 5 dimes in your pocket (basically, ¢50). Alright, so you have this 1/4 + 1/2 (the 1/2 is for the 50¢ which is half a dollar) so you know that you have 75¢, right? 75¢ is also what you would have had you 3 quarters, so we can write the following thing 1/4 + 1/2 = 3/4 Is it clear so far?

OpenStudy (anonymous):

yea i kind of understand it

OpenStudy (anonymous):

let's say you remove that quarter from your pocket. You know have only 5 dimes, which is worth 50¢. 50¢ is half a dollar. 50¢ is also the worth of 2 quarters New equation is : 1/2 = 2/4 Now, aside from the logic we used here to explain it, here's how you can work it out mathematically. 1/2 * 1 = 1/2 1 can be written in a ton of different fractions. it can be 2/2 (2 halves make a whole), 3/3, 4/4, 80/80, basically anything where the top and bottom part of the fraction is the same. So lemme rewrite this diffrently : 1/2* 2/2 = 2/4 ok, so, what does multiplying by 2/2 actually do for you? it's the same thing as writing this :